# The Race between Technology and Woman:

### v8

@@ A.2.5 Magnitude of the effects of technological change
The magnitude of the first terms in equations (A46)–(A49) depends on the substitution parameters in the production function and the income share of ICT capital relative to the income shares of other factors.
@@
The magnitude of the second to the last terms in equations (A46)–(A49) depends on the differences between the aggregate elasticities, which in turn depend on the differences between the substitution parameters in the production function and the income share of ICT capital relative to the income shares of other factors.
@@
The effects of technological progress embodied in ICT capital on the income shares of male and female skilled labor are positive, while those on the income shares of male and female unskilled labor are negative, if $\sigma_{fh}<0$, $\sigma_{mh}\simeq 0$, and $\sigma_{fu},\sigma_{mu}>0$. Thus, the magnitude of the effects are all proportional to the differences in the substitution parameters in the production function and the income share of ICT capital relative to the income shares of other factors.
@@
The magnitude of the effect is proportional to the differences between the aggregate elasticities (A50), (A51), and (A54). Thus, the magnitude of the effects on the four wage gaps is proportional to the differences in the substitution parameters in the production function and the income share of ICT capital relative to the income shares of other factors.
A.3 Data description
A.3.1 Wages and hours worked
Suppose that there is no need to make an adjustment to wages and hours worked. Let checks denote unadjusted values. The wages for labor of type $f$ in sector $n$, country $j$, and year $t$ could be calculated as $\check{w}_{f,njt}=\sum_{a_{f}}\rho_{f,njt}^{a_{f}}\check{w}_{f,njt}^{a_{f}}$, where $\rho_{f,njt}^{a_{f}}$ is the share of total hours worked by group $a_{f}$ (i.e., $\rho_{f,njt}^{a_{f}}=\check{\ell}_{f,njt}^{a_{f}}/\sum_{a_{f}}\check{\ell}_{f,njt}^{a_{f}}$). The hours worked by labor of type $f$ in sector $n$, country $j$, and year $t$ could be calculated as ˇ$\check{\ell}_{f,njt}=\sum_{a_{f}}\check{\ell}_{f,njt}^{a_{f}}$. We adjust for changes in the age and education composition of the labor force by holding the share of each group constant when we calculate wages and by using the time-invariant efficiency units as weights when we calculate hours worked. Let $T_{j}$ denote the number of years observed for country $j$. The composition-adjusted wages for labor of type $f$ in sector $n$, country $j$, and year $t$ can be calculated as $w_{f,njt}=\sum_{a_{f}}\overline{\rho}_{f,nj}^{a_{f}}\check{w}_{f,njt}^{a_{f}}$, where $\overline{\rho}_{f,nj}^{a_{f}}$ is the sector- and country-specific mean of $\rho_{f,njt}^{a_{f}}$ (i.e., $\overline{\rho}_{f,nj}^{a_{f}}=\sum_{t=1}^{T_{j}}\rho_{f,njt}^{a_{f}}/T_{j}$). The composition-adjusted hours worked by labor of type $f$ in sector $n$, country $j$, and year $t$ can be calculated as $\ell_{f,njt}=\sum_{a_{f}}(\overline{w}_{f,nj}^{a_{f}}/\overline{w}_{f,nj}^{a_{f^{\prime}}})\check{\ell}_{f,njt}^{a_{f}}$, where the efficiency unit is measured by the sector- and country-specific mean of $\check{w}_{f,njt}^{a_{f}}$ (i.e., $\overline{w}_{f,nj}^{a_{f}}=\sum_{t=1}^{T_{j}}\check{w}_{f,njt}^{a_{f}}/T_{j}$) and normalized by $\overline{w}_{f,nj}^{a_{f^{\prime}}}$. We choose middle-aged high-skilled labor as the base group for skilled labor and middle-aged medium-skilled labor as the base group for unskilled labor. Our results do not depend on the choice of the base group.
A.3.2 Rental price of capital
where the interest rate is the internal rate of return:
and the interest rate as
where $cpi$ is the consumer price index. The data can be obtained from OECD.Stat.
Figure A1: ICT equipment, non-ICT equipment, and non-ICT structures
(a) Rental prices
(b) Quantities
Notes: The rental prices of ICT equipment, non-ICT equipment, and non-ICT structures are denoted by $r_{i}$, $r_{oe}$, and $r_{os}$, respectively, while the quantities of ICT equipment, non-ICT equipment, and non-ICT structures are denoted by $k_{i}$, $k_{oe}$, and $k_{os}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980. The 1980 values of $r_{i}$, $r_{oe}$, and $r_{os}$ are 1.14, 0.24, and 0.06, respectively. The 1980 values of $k_{i}$, $k_{oe}$, and $k_{os}$ are 31.14, 387.93, and 2000.84 billion U.S. dollars, respectively.
A.3.3 Non-ICT equipment
The trends in the rental prices of capital differ significantly between ICT equipment and non-ICT structures but do not differ significantly between non-ICT equipment and non-ICT structures (Figure A1a). The rental price of ICT equipment fell dramatically, but that of non-ICT equipment and structures remained almost unchanged. Meanwhile, the quantities of ICT equipment, non-ICT equipment, and non-ICT structures increased. However, the rate of increase in ICT equipment is far greater than that in non-ICT equipment and structures (Figure A1b). The rate of increase in non-ICT equipment is almost the same as that in non-ICT structures.

### v9

@@ E.5 Magnitude of the effects of technological change
The magnitude of the first terms in equations (A51)–(A54) depends on the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
@@
The magnitude of the second to the last terms in equations (A51)–(A54) depends on the differences between the aggregate elasticities, which in turn depend on the differences between the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
@@
The effects of technological progress embodied in ICCT capital on the income shares of male and female skilled labor are positive, while those on the income shares of male and female unskilled labor are negative, if $\sigma_{fh}<0$, $\sigma_{mh}\simeq 0$, and $\sigma_{fu},\sigma_{mu}>0$. Thus, the magnitude of the effects are all proportional to the differences in the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
@@
The magnitude of the effect is proportional to the differences between the aggregate elasticities (A55), (A56), and (A59). Thus, the magnitude of the effects on the four wage gaps is proportional to the differences in the substitution parameters in the production function and the income share of ICCT capital relative to the income shares of other factors.
Appendix F Additional Results
Table A2 report the changes attributable to the three types of technological change when labor is not perfectly mobile across sectors. Table A3 reports the effects of technological change embodied in ICCT capital on the relative wages and income shares of the four types of labor by country for the cases in which the supply of labor is perfectly inelastic and unit elastic. Table A4 compares the magnitude of the changes in relative wages and income shares attributable to ICCT equipment obtained by estimating and aggregating sectoral production functions with that obtained by assuming and estimating the aggregate production function.
Table A2: Effects of technological change under imperfect labor mobility, 1980–2005
$\Delta\ln q_{i}$ $\Delta\ln q_{o}$ $\Delta\ln A_{n}$
Relative wages
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 936 –0. 016 0. 072
(0. 056) (0. 001) (0. 005)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 062 –0. 001 –0. 004
(0. 006) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 240 0. 004 –0. 010
(0. 014) (0. 000) (0. 001)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 337 0. 022 –0. 104
(0. 058) (0. 001) (0. 005)
$\Delta\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 251 –0. 004 0. 014
(0. 015) (0. 000) (0. 001)
$\Delta\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 564 0. 009 –0. 040
(0. 013) (0. 000) (0. 001)
Labor shares
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 005 0. 000 0. 000
(0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 000 –0. 001
(0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 036 –0. 001 0. 002
(0. 002) (0. 000) (0. 000)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 0. 000 0. 001
(0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ 0. 001 0. 000 0. 000
(0. 004) (0. 000) (0. 000)
Notes: The changes in the levels of relative wages and labor shares attributable to technological change are computed using equations (A3) and (A4) of Proposition 5. Standard errors in parentheses are computed using bootstrap with 500 replications.
Table A3: Effects of technological change embodied in ICCT capital for each country
All AUS AUT CZE DEU DNK FIN GBR ITA JPN NLD USA
Relative wages, $1/\gamma=0$
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 886 –1. 351 –0. 565 –0. 770 –0. 879 –3. 325 –0. 264 –0. 938 –0. 552 –1. 318 –1. 301 –0. 770
(0. 052) (0. 082) (0. 033) (0. 048) (0. 053) (0. 239) (0. 013) (0. 057) (0. 036) (0. 099) (0. 092) (0. 043)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 074 –0. 140 –0. 038 –0. 070 –0. 058 –0. 205 –0. 025 –0. 085 –0. 043 –0. 089 –0. 081 –0. 081
(0. 006) (0. 010) (0. 003) (0. 006) (0. 005) (0. 017) (0. 003) (0. 007) (0. 004) (0. 009) (0. 007) (0. 007)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 262 0. 412 0. 226 0. 306 0. 216 1. 105 0. 102 0. 225 0. 223 0. 188 0. 419 0. 176
(0. 014) (0. 022) (0. 014) (0. 016) (0. 011) (0. 075) (0. 005) (0. 012) (0. 015) (0. 008) (0. 022) (0. 008)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 283 1. 822 1. 037 1. 359 1. 239 5. 308 0. 303 1. 446 0. 829 1. 665 2. 240 0. 934
(0. 053) (0. 076) (0. 041) (0. 062) (0. 056) (0. 274) (0. 009) (0. 064) (0. 030) (0. 108) (0. 121) (0. 037)
Relative wages, $1/\gamma=1$
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 486 –0. 737 –0. 315 –0. 420 –0. 470 –1. 599 –0. 155 –0. 503 –0. 299 –0. 660 –0. 655 –0. 431
(0. 036) (0. 045) (0. 019) (0. 027) (0. 028) (0. 110) (0. 006) (0. 030) (0. 017) (0. 047) (0. 044) (0. 024)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 096 –0. 177 –0. 050 –0. 085 –0. 074 –0. 238 –0. 033 –0. 109 –0. 053 –0. 102 –0. 100 –0. 108
(0. 010) (0. 021) (0. 007) (0. 011) (0. 010) (0. 032) (0. 006) (0. 015) (0. 008) (0. 014) (0. 013) (0. 014)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 264 0. 422 0. 230 0. 303 0. 212 0. 992 0. 105 0. 229 0. 206 0. 177 0. 400 0. 189
(0. 013) (0. 020) (0. 012) (0. 017) (0. 011) (0. 058) (0. 009) (0. 012) (0. 011) (0. 012) (0. 022) (0. 011)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 0. 767 1. 092 0. 634 0. 807 0. 717 2. 853 0. 196 0. 844 0. 507 0. 876 1. 226 0. 567
(0. 039) (0. 044) (0. 025) (0. 038) (0. 032) (0. 131) (0. 006) (0. 037) (0. 017) (0. 054) (0. 059) (0. 023)
Labor shares, $1/\gamma=0$
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 074 –0. 025 –0. 027 –0. 024 –0. 107 –0. 013 –0. 037 –0. 032 –0. 032 –0. 050 –0. 036
(0. 002) (0. 004) (0. 002) (0. 002) (0. 002) (0. 007) (0. 001) (0. 002) (0. 002) (0. 002) (0. 003) (0. 002)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 –0. 011 –0. 006 –0. 007 –0. 005 –0. 027 –0. 005 –0. 006 –0. 005 –0. 004 –0. 004 –0. 008
Labor shares, $1/\gamma=1$
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 014 0. 014 0. 010 0. 010 0. 011 0. 028 0. 004 0. 009 0. 008 0. 012 0. 019 0. 010
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 002) (0. 000) (0. 000) (0. 000) (0. 001) (0. 001) (0. 001)
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 000) (0. 000) (0. 000) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 070 –0. 128 –0. 046 –0. 046 –0. 042 –0. 173 –0. 024 –0. 066 –0. 052 –0. 051 –0. 083 –0. 067
(0. 005) (0. 009) (0. 004) (0. 004) (0. 004) (0. 016) (0. 003) (0. 006) (0. 005) (0. 005) (0. 008) (0. 006)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 007 –0. 009 –0. 005 –0. 008 –0. 004 –0. 019 –0. 007 –0. 004 –0. 003 –0. 003 0. 001 –0. 010
Technological change embodied in ICCT capital
Notes: Country names are abbreviated as follows: AUS, Australia; AUT, Austria; CZE, the Czech Republic; DEU, Germany; DNK, Denmark; FIN, Finland; GBR, the United Kingdom; ITA, Italy; JPN, Japan; NLD, the Netherlands; and USA, the United States. The changes in the levels of relative wages and labor shares attributable to technological change are computed using equations (A1) and (A2) of Proposition 4. The first column reproduces the first and third columns of Table 7. The second to last columns report the effects evaluated at the sample means over time for each country. Standard errors in parentheses are computed using bootstrap with 500 replications.
Table A4: Results from sectoral and aggregate production functions
$\Delta\ln k_{i}$ $\Delta\ln q_{i}$
Sectoral Aggregate Sectoral Aggregate
Relative wages Relative wages
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 884 –1. 109 $\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 886 –1. 015
(0. 070) (0. 097) (0. 052) (0. 061)
$\Delta\ln\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 070 –0. 075 $\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 074 –0. 073
(0. 005) (0. 006) (0. 006) (0. 005)
$\Delta\ln\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 222 0. 236 $\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 262 0. 254
(0. 014) (0. 017) (0. 014) (0. 015)
$\Delta\ln\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 036 1. 270 $\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 283 1. 436
(0. 069) (0. 096) (0. 053) (0. 063)
$\Delta\ln\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 228 –0. 276 $\Delta\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 249 –0. 276
(0. 017) (0. 022) (0. 014) (0. 015)
$\Delta\ln\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 454 0. 531 $\Delta\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 564 0. 602
(0. 024) (0. 032) (0. 013) (0. 014)
Labor shares Labor shares
$\Delta\ln\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 092 0. 101 $\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 006
(0. 015) (0. 018) (0. 001) (0. 001)
$\Delta\ln\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 976 1. 211 $\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 019
(0. 071) (0. 098) (0. 001) (0. 001)
$\Delta\ln\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 130 –0. 135 $\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 037
(0. 004) (0. 003) (0. 002) (0. 002)
$\Delta\ln\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 060 –0. 059 $\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 –0. 007
(0. 005) (0. 005) (0. 001) (0. 001)
$\Delta\ln\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 003 0. 012 $\Delta\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 001 0. 006
(0. 007) (0. 009) (0. 004) (0. 004)
Notes: The figures in the upper and lower halves of the first two columns are computed using equation (18) of Proposition 2 and equation (15) of Proposition 1, while the figures in the upper and lower halves of the last two columns are computed using equations (21) and (22) of Proposition 3. The figures in the first column of the left panel are reproduced from the fifth columns of Tables 5 and 6, while those in the first column of the right panel are reproduced from the first column of Table 7. The figures in the second columns of the left and right panels are computed using the estimates of the aggregate production function parameters reported in the last row of Table 2.